OpenStudy (michaelbp):

What is the equation of a line that passes through (7, 8) and has a slope of -3?

OpenStudy (ovidia):

not gonna lie this one is kinda tricky bro

OpenStudy (michaelbp):

i knoww

OpenStudy (ovidia):

ight well let me see whats up then hmmmm?

OpenStudy (isaidavila):

Remember point slope equation (y-yo)=m(x-xo) You got (xo,yo)=(7,8) and m=-3

OpenStudy (ovidia):

its A. because B slope is 3 C slope is 1/3 and D slope is -1/3 A is the only -3

OpenStudy (ovidia):

@michaelbp

OpenStudy (mhchen):

Any ideas on how to start? There's 2 formulas you can use. Point-slope and slope-intercept form. I recommend point slope. You've studied that in class right?